# GPR: Towards a Generative Pre-trained One-Model Paradigm for Large-Scale Advertising Recommendation

### v2

@@ 1. Introduction
(su2009survey; huang2013learning; cheng2016wide; zhou2018deep; zhang2024wukong; zhu2025rankmixer), as shown in Fig. 1(a). However, such pipelines suffer from inconsistent optimization objectives across stages: the retrieval stage emphasizes coverage optimization, while the final ranking stage focuses on business outcome prediction. This misalignment prevents the system from achieving global optimality. Moreover, the limited representational capacity of early retrieval models often leads to the premature elimination of potentially high-quality candidates, restricting subsequent ranking models from fully exploiting deep feature interactions and resulting in an information bottleneck. In addition, the cascading architecture relies on highly complex engineering implementations, requiring substantial resources to maintain cross-stage consistency, which in turn hinders rapid algorithmic iteration and system scalability.
In recent years, with the rapid advancement of large language models (LLMs)  (vaswani2017attention; achiam2023gpt; guo2025deepseek), particularly their breakthroughs in generative capabilities (zhai2024actions; rajput2023recommender), the research paradigm of recommendation systems has been continuously evolving. HSTU (zhai2024actions) introduced generative modeling into the traditional cascading architecture, as shown in Fig. 1(b), further driving the transition from the multi-stage pipeline to a unified generative recommendation framework (zhou2025onerec; han2025mtgr). Such models can directly generate optimal recommendations based on a holistic understanding of user interests and semantic context, thereby ensuring the consistency of optimization objectives. However, due to the inherent complexity of advertising recommendation tasks, existing generative models still face multiple challenges in real-world industrial applications:
@@
2) Efficiency–Flexibility Trade-off: In industrial-scale advertising recommendation, the model must not only support efficient training for large-scale data updates but also provide flexible decoding capabilities to process ultra-long user behavior sequences in real time and match ads under multiple constraints (e.g., targeting, bidding, budget). Decoder-only architectures (zhai2024actions) enable efficient user-level training but suffer from limited decoding flexibility, whereas encoder–decoder architectures (zhou2025onerec) allow flexible inference but incur prohibitively high training costs due to pointwise loss objectives, making them unsuitable for real-time advertising systems. Balancing efficient training with flexible inference thus remains a key challenge.
@@ 2.1. Input Schema and Processing
Furthermore, as demonstrated by previous GR models (rajput2023recommender; zhou2025onerec), quantizing item embeddings into semantic IDs can extract crucial information in items, while also generate discrete representation of items, better aligning with generative model paradigms. Therefore, we convert all contents in O-Token and items in I-Token into discrete semantic IDs. RQ-VAE (rq-vae) and RQ-Kmeans (rq-kmeans) are two popular methods. However, both of them suffer from issues such as ”codebook collapse” and ”insufficient robustness of the latent space”, leading to low utilization of the semantic space and limited representational capacity. To address these issues, we propose a new quantization model named RQ-Kmeans+, which consists of an encoder, residual codebooks and a decoder.
@@ 2.2. Heterogeneous Hierarchical Decoder
The HSD module is a primary decoder, which stacks HSTU blocks (zhai2024actions) and takes the unified token sequence as input to understand users’ actions and generate intent embeddings. Building upon the foundational HSTU block, we introduce several critical enhancements to form the block utilized in the HSD, including the Hybrid Attention mechanism, Token-Aware Normalization, and an improved Feed Forward Network (FFN), as shown in Fig.  2 (b).
Unlike standard attention mechanisms, the HSD Attention mechanism incorporates an extra embedding, $U$, which adaptively modulate the attention weights, $W$. This modulation enables the HSD to focus more effectively on relevant user behaviors while actively attenuating less informative interactions (zhai2024actions). Moreover, given that the U-Token, O-Token, and E-Token in our input schema collectively function as a prompt for item prediction, applying a vanilla causal mask within this prompt region becomes unnecessarily restrictive. To address this limitation, we propose the Hybrid Attention mask, $M^{hybrid}$. Within the prefix block, tokens visible with each other freely using bi-directional attention. This design enables the model to fully exploit the contextual interplay among the prompt tokens, thereby constructing a more comprehensive context prior to making a prediction. The Hybrid Attention can be expressed as:
@@
In addition, considering that different types of tokens possess distinct characteristics, the HSD module assigns independent normalization layers and feed-forward networks (FFNs) to each token type, projecting them into their own semantic subspaces to fully capture the semantic diversity of heterogeneous sequences. Meanwhile, we introduce the Mixture-of-Recursions (MoR) mechanism  (bae2025mixture), which increases the model’s depth and reasoning capacity without adding extra parameters. To further enhance reasoning ability, the model incorporates external knowledge from a fine-tuned Large Language Model (LLM), where the LLM generates a textual “thought process” about users’ potential interests, which is then tokenized and integrated into the intent embeddings to strengthen semantic understanding and reasoning capability.
@@
Moreover, inspired by recent research on reasoning in LLM (madaan2023self; wu2024large), we extend PTD by integrating a refining module to further strengthen its cognitive abilities and generative capacities. As shown in Fig.  2 (c), the refining module is designed upon the diffusion paradigm (rombach2022high), which consists of a noise generator and a reverse process modeled as a Markov chain. In the reverse process, the noise is iteratively removed by a conditional denoising module with a transformer architecture. Prefix thinking tokens are aggregated using Sum_Pooling and serve as the condition for the denoising module. Consequently, the refining module refines the initial reasoning results, which are subsequently passed to downstream modules.
@@ 3.1. Pre-training with Multi-Token Prediction
Items (ads) are encoded as $L$ coarse-to-fine semantic codes obtained via residual vector quantization, which provides a hierarchical, compact representation of item semantics. As a baseline, next-token prediction (NTP) conditions on the user history and context to predict the subsequent code. While effective for single-path dynamics, NTP implicitly assumes a dominant interest trajectory and tends to average over parallel interests, limiting coverage in advertising scenarios. To capture multiple concurrent interests, as shown in Fig.  4 (a), we employ MTP (gloeckle2024better) by extending decoder with $N$ parallel heads (default $N=4$). Each head independently predicts a complete $L$-level code path for one interest dimension, using the same backbone states but separate projection layers. This design enables concurrent modeling without mutual interference and preserves level-wise legality via masked decoding on each head.
@@ 4.1. Multimodal Tokenization Performance
Experiment setting. A large-scale corpus from Tencent’s advertising platform is curated, covering ads and organic media (short videos, social feeds, news). The corpus comprises heterogeneous multimodal signals—textual metadata (titles, tags, descriptions) and visual content (thumbnails and sampled frames) —aligned at the item and session level. Near-duplicate samples are filtered to reduce redundancy and category distributions are balanced to limit sampling bias. The corpus is partitioned into 80% training and 20% test. RQ-KMeans+ is compared against two established tokenizers: RQ-VAE (rq-vae) (standard discrete tokenizer prone to dead codes) and RQ-KMeans (rq-kmeans) (K-means initialization without VAE-style refinement). We introduce three metrics to assess code quality. Specifically, Collision Rate measures the fraction of items mapped to an identical code and lower values are better. Code Usage Rate at level 1 (CURL1) measures the proportion of active codes and higher values are better. Path Average Similarity (PAS) measures the mean embedding similarity among items that share a code and higher PAS indicate more coherent collisions.
@@ 4.2. User Behavior Modeling Performance
Experiment setting. We use a large-scale Tencent advertising dataset (user interactions with ads and organic content). Training uses one year of anonymized user interactions and validation uses the next calendar day. Histories are serialized in strict time order with the unified four-token schema (U/O/E/I). The same tokenizer (RQ-KMeans+) and feature pipelines are shared by all models. Baselines are HSTU (zhai2024actions) (decoder-only) and OneRec (zhou2025onerec2) (encoder–decoder). The task is to predict the next interacted item, and each model is required to emit top 100 results from the full catalog snapshot, which contains millions of items. HitRate@100 is defined by this retrieval protocol: a hit is recorded if the ground-truth ad appears among the top-100 candidates. This metric avoids overly strict code equality and evaluates whether the generated semantic code localizes the correct item in embedding space.
@@ 4.3. Business Alignment Performance
Experiment setting. This study evaluates whether the joint training strategy—MTP pretraining, eCPM-aware fine-tuning, and HEPO— improves monetization alignment. The dataset follows Section 4.2: one year for training and the subsequent day for validation. In addition to sequence signals, labels include eCPM and action types (impression, click, conversion) to reflect value hierarchy. Methods compared are: MTP, a baseline multi-token prediction model trained with uniform head/position weights under likelihood only; MTP+VAFT, a business-aligned variant that reweights the MTP loss at each position using a value-aware factor; MTP+DPO (rafailov2023direct), a preference-based fine-tuning that learns from pairwise orderings constructed under predicted rewards; and MTP+HEPO. We evaluate with nDCG (normalized Discounted Cumulative Gain), which measures ranked-list quality with higher weight on the top, and OPR (Ordered Pair Ratio), which measures the fraction of item pairs ordered correctly. Both metrics are computed under the eCPM-oriented ranking.
@@ 4.4. Online Performance
Experiment setting. We evaluate and depoly GPR on Weixin Channels, a production advertising platform with hundreds of millions of active users and tens of millions of dynamic ads. The comparison baseline is a mature multi-stage cascade with multiple retrieval methods and customized strategies. Primary online metrics include GMV, Costs, and CTCVR, with GMV as the primary KPI because it directly reflects business return. We conduct five sequential online A/B evaluations of the GPR stack. Each iteration undergoes rigorous validation and produces stable, statistically significant incremental improvements over the preceding version on the primary KPIs.
@@
The stratified analysis based on first launch (v0.1) in Tab.  5 shows consistent gains across users and ad segements. UG1 and UG2 are low-activity users and show strong gains, for example UG1 reaches GMV +3.56%, CTR +2.51%, CVR +0.82%, and CTCVR +3.72%. Mid-activity groups UG3 and UG4 also improve across engagement and efficiency, with UG3 showing the largest CTCVR lift at +4.63%. UG5 denotes high-activity users and exhibits smaller CTR and CVR changes yet still delivers GMV +3.68%, indicating better allocation toward higher-value ads even for heavy users. For inventory, newly launched ads ($\leq$3 days) outperform established ads ads (¿3 days ) with GMV +2.97% vs. +1.65% and CTCVR +4.02% vs. +2.78%. This pattern indicates stronger cold-start handling while preserving gains for mature inventory.
@@ 5. Related Work
LLMs as Generative Rankers. In the last two years, researchers have begun exploring the application of LLMs in recommender systems, primarily leveraging their ability to generate the next item recommendation (wu2024survey). In these efforts, user behavior sequences are typically transformed into text sequences, and the LLMs are tasked with generating the next item based on this textual input (zhu2024collaborative; lin2024rella; li2024leadre; zheng2024adapting; harte2023leveraging; zhang2023chatgpt). However, a core limitation of this approach is the fixed vocabulary of LLMs, which hinders their ability to adapt effectively to the dynamically changing and large-scale item sets characteristic of modern advertising recommendation scenarios.
@@
Generative Recommendation Models. Recently, driven by the success of LLMs, there has been a significant shift towards generative recommendation (rajput2023recommender; chen2024hllm; zhai2024actions; liu2024kuaiformer; badrinath2025pinrec; yang2024unifying; qiu2025one; huang2025towards; wang2025scaling; jiang2025large). TIGER (rajput2023recommender) is a generative recommendation model that leverages semantic ids, using a sequence-to-sequence framework to generate such semantic ids for recommendations. HLLM (chen2024hllm) leverages the pre-trained capabilities of LLMs and models item representations and user interests respectively via a two-layer architecture. HSTU (zhai2024actions) adopts a decoder-only architecture to process ultra-long user histories and generate item recommendations. These work have demonstrated the potential of generative models to unify different recommendation tasks and handle long user sequences. MTGR (han2025mtgr) adopts the HSTU architecture while retaining the original deep learning recommendation model (DLRM) features. COBRA (yang2025sparse) integrates sparse semantic IDs and dense vectors through a cascading process. GPR builds on this generative trend but introduces a novel HHD architecture. Unlike previous approaches which entangle user representation with item generation, GPR decouples these two tasks via a hierarchical structure and propose a novel ”understanding-thinking-refining-generation” paradigm.
@@
End-to-End Recommendation Frameworks. Recent studies have attempted to construct more unified, end-to-end recommendation systems, such as OneRec(zhou2025onerec). OneRec unifies retrieve-and-rank via a encoder-decoder architecture and a preference alignment algorithm (e.g., DPO) in a video recommendation system. By contrast, GPR is the first end-to-end generative solution successfully deployed in large-scale advertising system, which faces unique challenges including behavior heterogeneity and sparsity, multi-objective optimization, precise value prediction, etc.
@@
RL in Recommender Systems. In traditional recommender systems, reinforcement learning (RL) is used to optimize slate/page decisions and long-horizon value, rather than just next-item accuracy. Seq2Slate (bello2018seq2slate) formulates re-ranking as autoregressive slate generation and trains via policy gradients, enabling the model to assign credit to the entire slate (instead of treating items independently) and capture cross-position effects that pointwise losses miss. SlateQ (ie2019reinforcement) shows that, under mild user-choice assumptions, the long-term value of a slate can be decomposed into itemwise terms—making TD/Q-learning tractable for page-wise optimization and enabling production-scale experiments. At an extremely large scale, YouTube’s system introduces Top-K off-policy correction to learn the recommendation policy from logged implicit feedback, while incorporating a learned model of the behavior policy to adjust data biases. Ad-mixed feeds further highlight the need for RL, as the system must jointly balance platform revenue, advertiser ROI, and user experience at the slate level: DEAR (zhao2021dear) addresses ad insertion as a joint decision (whether to insert an ad, which ad to show, and where to place it) using a DQN that optimizes long-run value under the delayed/mixed rewards typical of advertising. To reduce online risk, teams often start with offline RL; a common choice is Conservative Q-Learning (CQL) (kumar2020conservative), which penalizes out-of-distribution actions to prevent the learned value function from overestimating unseen choices before online fine-tuning.
@@
RL for Generative Recommendation. In generative recommendation, RL is used to align sequence-producing policies with user and business preferences. GeMS (deffayet2023generative) learns a variational latent space for slates, enabling an RL agent (e.g., SAC) to act in this continuous space before decoding full lists—sidestepping combinatorial action spaces while still optimizing sequence-level rewards. When hand-crafted rewards are brittle, PrefRec (xue2023prefrec) first learns a reward model from human/trajectory preferences, then optimizes policies against this model to improve long-term engagement. Direct Preference Optimization (DPO)-style objectives (rafailov2023direct) adapt this idea to sequence models (e.g., Softmax-DPO; DPO4Rec), turning pairwise preferences into stable training losses that can replace or complement RL during post-training (e.g., for aligning generative policies with user preferences). Recent industrial systems combine these elements end-to-end: OneRec-V1 (zhou2025onerec) unifies retrieve-and-rank with a generator and iterative preference alignment (via DPO), reporting a +1.6% online watch-time lift; OneRec-V2 (zhou2025onerec2) shifts to a lazy decoder-only stack and adds RL on real user feedback (with duration-aware reward shaping and adaptive ratio clipping) to optimize long-term engagement.

### v3

@@ 1. Introduction
(Su and Khoshgoftaar, 2009; Huang et al., 2013; Cheng et al., 2016; Zhou et al., 2018; Zhang et al., 2024; Zhu et al., 2025), as shown in Fig. 1(a). However, such pipelines suffer from inconsistent optimization objectives across stages: the retrieval stage emphasizes coverage optimization, while the final ranking stage focuses on business outcome prediction. This misalignment prevents the system from achieving global optimality. Moreover, the limited representational capacity of early retrieval models often leads to the premature elimination of potentially high-quality candidates, restricting subsequent ranking models from fully exploiting deep feature interactions and resulting in an information bottleneck. In addition, the cascading architecture relies on highly complex engineering implementations, requiring substantial resources to maintain cross-stage consistency, which in turn hinders rapid algorithmic iteration and system scalability.
In recent years, with the rapid advancement of large language models (LLMs)  (Vaswani et al., 2017; Achiam et al., 2023; Guo et al., 2025), particularly their breakthroughs in generative capabilities (Zhai et al., 2024; Rajput et al., 2023), the research paradigm of recommendation systems has been continuously evolving. HSTU (Zhai et al., 2024) introduced generative modeling into the traditional cascading architecture, as shown in Fig. 1(b), further driving the transition from the multi-stage pipeline to a unified generative recommendation framework (Zhou et al., 2025a; Han et al., 2025). Such models can directly generate optimal recommendations based on a holistic understanding of user interests and semantic context, thereby ensuring the consistency of optimization objectives. However, due to the inherent complexity of advertising recommendation tasks, existing generative models still face multiple challenges in real-world industrial applications:
@@
2) Efficiency–Flexibility Trade-off: In industrial-scale advertising recommendation, the model must not only support efficient training for large-scale data updates but also provide flexible decoding capabilities to process ultra-long user behavior sequences in real time and match ads under multiple constraints (e.g., targeting, bidding, budget). Decoder-only architectures (Zhai et al., 2024) enable efficient user-level training but suffer from limited decoding flexibility, whereas encoder–decoder architectures (Zhou et al., 2025a) allow flexible inference but incur prohibitively high training costs due to pointwise loss objectives, making them unsuitable for real-time advertising systems. Balancing efficient training with flexible inference thus remains a key challenge.
@@ 2.1. Input Schema and Processing
Furthermore, as demonstrated by previous GR models (Rajput et al., 2023; Zhou et al., 2025a), quantizing item embeddings into semantic IDs can extract crucial information in items, while also generate discrete representation of items, better aligning with generative model paradigms. Therefore, we convert all contents in O-Token and items in I-Token into discrete semantic IDs. RQ-VAE (Lee et al., 2022) and RQ-Kmeans (Luo et al., 2024) are two popular methods. However, both of them suffer from issues such as ”codebook collapse” and ”insufficient robustness of the latent space”, leading to low utilization of the semantic space and limited representational capacity. To address these issues, we propose a new quantization model named RQ-Kmeans+, which consists of an encoder, residual codebooks and a decoder.
@@ 2.2. Heterogeneous Hierarchical Decoder
The HSD module is a primary decoder, which stacks HSTU blocks (Zhai et al., 2024) and takes the unified token sequence as input to understand users’ actions and generate intent embeddings. Building upon the foundational HSTU block, we introduce several critical enhancements to form the block utilized in the HSD, including the Hybrid Attention mechanism, Token-Aware Normalization, and an improved Feed Forward Network (FFN), as shown in Fig.  2 (b).
Unlike standard attention mechanisms, the HSD Attention mechanism incorporates an extra embedding, $U$, which adaptively modulate the attention weights, $W$. This modulation enables the HSD to focus more effectively on relevant user behaviors while actively attenuating less informative interactions (Zhai et al., 2024). Moreover, given that the U-Token, O-Token, and E-Token in our input schema collectively function as a prompt for item prediction, applying a vanilla causal mask within this prompt region becomes unnecessarily restrictive. To address this limitation, we propose the Hybrid Attention mask, $M^{hybrid}$. Within the prefix block, tokens visible with each other freely using bi-directional attention. This design enables the model to fully exploit the contextual interplay among the prompt tokens, thereby constructing a more comprehensive context prior to making a prediction. The Hybrid Attention can be expressed as:
@@
In addition, considering that different types of tokens possess distinct characteristics, the HSD module assigns independent normalization layers and feed-forward networks (FFNs) to each token type, projecting them into their own semantic subspaces to fully capture the semantic diversity of heterogeneous sequences. Meanwhile, we introduce the Mixture-of-Recursions (MoR) mechanism  (Bae et al., 2025), which increases the model’s depth and reasoning capacity without adding extra parameters. To further enhance reasoning ability, the model incorporates external knowledge from a fine-tuned Large Language Model (LLM), where the LLM generates a textual “thought process” about users’ potential interests, which is then tokenized and integrated into the intent embeddings to strengthen semantic understanding and reasoning capability.
@@
Moreover, inspired by recent research on reasoning in LLM (Madaan et al., 2023; Wu et al., 2024b), we extend PTD by integrating a refining module to further strengthen its cognitive abilities and generative capacities. As shown in Fig.  2 (c), the refining module is designed upon the diffusion paradigm (Rombach et al., 2022), which consists of a noise generator and a reverse process modeled as a Markov chain. In the reverse process, the noise is iteratively removed by a conditional denoising module with a transformer architecture. Prefix thinking tokens are aggregated using Sum_Pooling and serve as the condition for the denoising module. Consequently, the refining module refines the initial reasoning results, which are subsequently passed to downstream modules.
@@ 3.1. Pre-training with Multi-Token Prediction
Items (ads) are encoded as $L$ coarse-to-fine semantic codes obtained via residual vector quantization, which provides a hierarchical, compact representation of item semantics. As a baseline, next-token prediction (NTP) conditions on the user history and context to predict the subsequent code. While effective for single-path dynamics, NTP implicitly assumes a dominant interest trajectory and tends to average over parallel interests, limiting coverage in advertising scenarios. To capture multiple concurrent interests, as shown in Fig.  4 (a), we employ MTP (Gloeckle et al., 2024) by extending decoder with $N$ parallel heads (default $N=4$). Each head independently predicts a complete $L$-level code path for one interest dimension, using the same backbone states but separate projection layers. This design enables concurrent modeling without mutual interference and preserves level-wise legality via masked decoding on each head.
@@ 4.1. Multimodal Tokenization Performance
Experiment setting. A large-scale corpus from Tencent’s advertising platform is curated, covering ads and organic media (short videos, social feeds, news). The corpus comprises heterogeneous multimodal signals—textual metadata (titles, tags, descriptions) and visual content (thumbnails and sampled frames) —aligned at the item and session level. Near-duplicate samples are filtered to reduce redundancy and category distributions are balanced to limit sampling bias. The corpus is partitioned into 80% training and 20% test. RQ-KMeans+ is compared against two established tokenizers: RQ-VAE (Lee et al., 2022) (standard discrete tokenizer prone to dead codes) and RQ-KMeans (Luo et al., 2024) (K-means initialization without VAE-style refinement). We introduce three metrics to assess code quality. Specifically, Collision Rate measures the fraction of items mapped to an identical code and lower values are better. Code Usage Rate at level 1 (CURL1) measures the proportion of active codes and higher values are better. Path Average Similarity (PAS) measures the mean embedding similarity among items that share a code and higher PAS indicate more coherent collisions.
@@ 4.2. User Behavior Modeling Performance
Experiment setting. We use a large-scale Tencent advertising dataset (user interactions with ads and organic content). Training uses one year of anonymized user interactions and validation uses the next calendar day. Histories are serialized in strict time order with the unified four-token schema (U/O/E/I). The same tokenizer (RQ-KMeans+) and feature pipelines are shared by all models. Baselines are HSTU (Zhai et al., 2024) (decoder-only) and OneRec (Zhou et al., 2025b) (encoder–decoder). The task is to predict the next interacted item, and each model is required to emit top 100 results from the full catalog snapshot, which contains millions of items. HitRate@100 is defined by this retrieval protocol: a hit is recorded if the ground-truth ad appears among the top-100 candidates. This metric avoids overly strict code equality and evaluates whether the generated semantic code localizes the correct item in embedding space.
@@ 4.3. Business Alignment Performance
Experiment setting. This study evaluates whether the joint training strategy—MTP pretraining, eCPM-aware fine-tuning, and HEPO— improves monetization alignment. The dataset follows Section 4.2: one year for training and the subsequent day for validation. In addition to sequence signals, labels include eCPM and action types (impression, click, conversion) to reflect value hierarchy. Methods compared are: MTP, a baseline multi-token prediction model trained with uniform head/position weights under likelihood only; MTP+VAFT, a business-aligned variant that reweights the MTP loss at each position using a value-aware factor; MTP+DPO (Rafailov et al., 2023), a preference-based fine-tuning that learns from pairwise orderings constructed under predicted rewards; and MTP+HEPO. We evaluate with nDCG (normalized Discounted Cumulative Gain), which measures ranked-list quality with higher weight on the top, and OPR (Ordered Pair Ratio), which measures the fraction of item pairs ordered correctly. Both metrics are computed under the eCPM-oriented ranking.
@@ 4.4. Online Performance
Experiment setting. We evaluate and deploy GPR on Weixin Channels, a production advertising platform with hundreds of millions of active users and tens of millions of dynamic ads. The comparison baseline is a mature multi-stage cascade with multiple retrieval methods and customized strategies. Primary online metrics include GMV, Costs, and CTCVR, with GMV as the primary KPI because it directly reflects business return. We conduct five sequential online A/B evaluations of the GPR stack. Each iteration undergoes rigorous validation and produces stable, statistically significant incremental improvements over the preceding version on the primary KPIs.
@@
The stratified analysis based on first launch (v0.1) in Tab.  5 shows consistent gains across users and ad segments. UG1 and UG2 are low-activity users and show strong gains, for example UG1 reaches GMV +3.56%, CTR +2.51%, CVR +0.82%, and CTCVR +3.72%. Mid-activity groups UG3 and UG4 also improve across engagement and efficiency, with UG3 showing the largest CTCVR lift at +4.63%. UG5 denotes high-activity users and exhibits smaller CTR and CVR changes yet still delivers GMV +3.68%, indicating better allocation toward higher-value ads even for heavy users. For inventory, newly launched ads ($\leq$3 days) outperform established ads (¿3 days ) with GMV +2.97% vs. +1.65% and CTCVR +4.02% vs. +2.78%. This pattern indicates stronger cold-start handling while preserving gains for mature inventory.
@@ 5. Related Work
LLMs as Generative Rankers. In the last two years, researchers have begun exploring the application of LLMs in recommender systems, primarily leveraging their ability to generate the next item recommendation (Wu et al., 2024a). In these efforts, user behavior sequences are typically transformed into text sequences, and the LLMs are tasked with generating the next item based on this textual input (Zhu et al., 2024; Lin et al., 2024; Li et al., 2024; Zheng et al., 2024; Harte et al., 2023; Zhang et al., 2023). However, a core limitation of this approach is the fixed vocabulary of LLMs, which hinders their ability to adapt effectively to the dynamically changing and large-scale item sets characteristic of modern advertising recommendation scenarios.
@@
Generative Recommendation Models. Recently, driven by the success of LLMs, there has been a significant shift towards generative recommendation (Rajput et al., 2023; Chen et al., 2024; Zhai et al., 2024; Liu et al., 2024; Badrinath et al., 2025; Yang et al., 2024; Qiu et al., 2025; Huang et al., 2025; Wang et al., 2025; Jiang et al., 2025). TIGER (Rajput et al., 2023) is a generative recommendation model that leverages semantic ids, using a sequence-to-sequence framework to generate such semantic ids for recommendations. HLLM (Chen et al., 2024) leverages the pre-trained capabilities of LLMs and models item representations and user interests respectively via a two-layer architecture. HSTU (Zhai et al., 2024) adopts a decoder-only architecture to process ultra-long user histories and generate item recommendations. These work have demonstrated the potential of generative models to unify different recommendation tasks and handle long user sequences. MTGR (Han et al., 2025) adopts the HSTU architecture while retaining the original deep learning recommendation model (DLRM) features. COBRA (Yang et al., 2025) integrates sparse semantic IDs and dense vectors through a cascading process. GPR builds on this generative trend but introduces a novel HHD architecture. Unlike previous approaches which entangle user representation with item generation, GPR decouples these two tasks via a hierarchical structure and propose a novel ”understanding-thinking-refining-generation” paradigm.
@@
End-to-End Recommendation Frameworks. Recent studies have attempted to construct more unified, end-to-end recommendation systems, such as OneRec(Zhou et al., 2025a). OneRec unifies retrieve-and-rank via a encoder-decoder architecture and a preference alignment algorithm (e.g., DPO) in a video recommendation system. By contrast, GPR is the first end-to-end generative solution successfully deployed in large-scale advertising system, which faces unique challenges including behavior heterogeneity and sparsity, multi-objective optimization, precise value prediction, etc.
@@
RL in Recommender Systems. In traditional recommender systems, reinforcement learning (RL) is used to optimize slate/page decisions and long-horizon value, rather than just next-item accuracy. Seq2Slate (Bello et al., 2018) formulates re-ranking as autoregressive slate generation and trains via policy gradients, enabling the model to assign credit to the entire slate (instead of treating items independently) and capture cross-position effects that pointwise losses miss. SlateQ (Ie et al., 2019) shows that, under mild user-choice assumptions, the long-term value of a slate can be decomposed into itemwise terms—making TD/Q-learning tractable for page-wise optimization and enabling production-scale experiments. At an extremely large scale, YouTube’s system introduces Top-K off-policy correction to learn the recommendation policy from logged implicit feedback, while incorporating a learned model of the behavior policy to adjust data biases. Ad-mixed feeds further highlight the need for RL, as the system must jointly balance platform revenue, advertiser ROI, and user experience at the slate level: DEAR (Zhao et al., 2021) addresses ad insertion as a joint decision (whether to insert an ad, which ad to show, and where to place it) using a DQN that optimizes long-run value under the delayed/mixed rewards typical of advertising. To reduce online risk, teams often start with offline RL; a common choice is Conservative Q-Learning (CQL) (Kumar et al., 2020), which penalizes out-of-distribution actions to prevent the learned value function from overestimating unseen choices before online fine-tuning.
@@
RL for Generative Recommendation. In generative recommendation, RL is used to align sequence-producing policies with user and business preferences. GeMS (Deffayet et al., 2023) learns a variational latent space for slates, enabling an RL agent (e.g., SAC) to act in this continuous space before decoding full lists—sidestepping combinatorial action spaces while still optimizing sequence-level rewards. When hand-crafted rewards are brittle, PrefRec (Xue et al., 2023) first learns a reward model from human/trajectory preferences, then optimizes policies against this model to improve long-term engagement. Direct Preference Optimization (DPO)-style objectives (Rafailov et al., 2023) adapt this idea to sequence models (e.g., Softmax-DPO; DPO4Rec), turning pairwise preferences into stable training losses that can replace or complement RL during post-training (e.g., for aligning generative policies with user preferences). Recent industrial systems combine these elements end-to-end: OneRec-V1 (Zhou et al., 2025a) unifies retrieve-and-rank with a generator and iterative preference alignment (via DPO), reporting a +1.6% online watch-time lift; OneRec-V2 (Zhou et al., 2025b) shifts to a lazy decoder-only stack and adds RL on real user feedback (with duration-aware reward shaping and adaptive ratio clipping) to optimize long-term engagement.